# Mirror Descent Policy Optimisation for Robust Constrained Markov Decision Processes

### v2

@@ Abstract
Safety is an essential requirement for reinforcement learning systems. The newly emerging framework of robust constrained Markov decision processes allows learning policies that satisfy long-term constraints while providing guarantees under epistemic uncertainty. This paper presents mirror descent policy optimisation for robust constrained Markov decision processes (RCMDPs), making use of policy gradient techniques to optimise both the policy (as a maximiser) and the transition kernel (as an adversarial minimiser) on the Lagrangian representing a constrained MDP. Our proposed algorithm obtains an $\tilde{\mathcal{O}}\left(1/T^{1/3}\right)$ convergence rate in the sample-based RCMDP setting. In addition to the RCMDP setting, the paper also contributes an algorithm for approximate gradient descent in the space of transition kernels, which is of independent interest for designing adversarial environments. Experiments confirm the benefits of mirror descent policy optimisation in constrained and unconstrained optimisation, and significant improvements are observed in robustness tests when compared to baseline policy optimisation algorithms.
Introduction
@@
Safe RL techniques have been primarily proposed within the constrained Markov decision process (CMDP) formalism (Altman,, 1999). In infinite MDPs, which are characterised by large or continuous state and action spaces, traditional solution techniques include linear programming and dynamic programming with Lagrangian relaxation. Following their empirical performance in high-dimensional control tasks (Lillicrap et al.,, 2015; Mnih et al.,, 2016), policy gradient algorithms have become a particularly popular option for safe RL in control applications (Achiam et al.,, 2017; Yang et al.,, 2020; Ding et al.,, 2020; Liu et al., 2021b, ; Xu et al.,, 2021; Paternain et al.,, 2023; Ying et al.,, 2023; Wu et al.,, 2024).
@@
The robustness to mismatches in training and test transition dynamics has been the main subject of interest in the robust MDP (RMDP) framework (Iyengar,, 2005; Nilim and Ghaoui,, 2005), which formulates robust policies by training them on the worst-case dynamics model within a plausible set called the uncertainty set. The framework has led to a variety of theoretically sound policy gradient approaches (Ho et al.,, 2021; Zhang et al.,, 2021; Kumar et al.,, 2023; Kuang et al.,, 2022; Wang et al.,, 2024; Li et al.,, 2023; Zhou et al.,, 2023; Wang et al.,, 2023). Recently, the approach has also branched out into the robust constrained MDP (RCMDP) framework (Russel et al.,, 2020, 2023; Wang et al.,, 2022; Bossens,, 2024; Zhang et al.,, 2024; Sun et al.,, 2024), which seeks to provide robustness guarantees to CMDPs.
@@
The framework of RCMDPs is newly emerging and hence the theoretical development lags behind compared to CMDPs and RMDPs. At present, a few challenges have hampered the progress of RCMDP policy optimisation. First, since the framework is based on a worst-case cost function and one or more additional worst-case constraint-cost functions, it is challenging to implement in a single simulator since each of the robust cost functions may turn out to be based on a different transition kernel. This has lead to practical objectives that do not meet the original objective (Russel et al.,, 2020) or which only work for a discrete set of transition kernels (Mankowitz et al.,, 2020). Second, the state-of-the-art CMDP techniques for policy optimisation and RMDP techniques for adversarially optimising the transition dynamics have not yet widely made their way into the RCMDP literature, which has lead to a lack of theoretical guarantees. In RMDPs, state-of-the-art results are obtained for mirror descent based policy gradient algorithm (Wang et al.,, 2024), which applies mirror descent in the policy space and mirror ascent in the transition kernel space. In CMDPs, the PMD-PD algorithm is a related mirror descent based policy gradient algorithm which provides $\mathcal{O}(\log(T)/T)$ convergence rate is obtained in the oracle-based setting and an $\tilde{\mathcal{O}}(1/T^{1/3})$ convergence rate in the sample-based setting (Liu et al., 2021b, ). We then pose the question: using mirror descent based policy optimisation, is it intrinsically more difficult to optimise RCMDPs, or is it possible to obtain similar convergence rates?
@@
With the aim of solving these challenges and providing an affirmative answer to the above question, we draw inspiration from the robust Lagrangian objective (Bossens,, 2024) and the success of mirror descent optimisation in RMDPs and CMDPs (Wang et al.,, 2024; Liu et al., 2021b, ). We show that it is indeed possible to obtain convergence rates similar to Sample-based PMD-PD in the sample-based RCMDP setting. For a theoretical analysis, we contribute Robust Sample-based PMD-PD, which extends the theory of Liu et al., 2021b and Wang et al., (2024) to the RCMDP setting, while focusing on the softmax parametrisation within the sample-based setting. Doing so, we achieve the $\tilde{\mathcal{O}}(1/T^{1/3})$ convergence rate as also obtained by (non-robust) Sample-based PMD-PD. For a practical implementation, we formulate MDPO-Robust-Lagrangian (MDPO-Robust-Lag for short), which introduces a robust Lagrangian to Mirror Descent Policy Optimisation (MDPO) (Tomar et al.,, 2022). We evaluate MDPO-Robust-Lagrangian empirically by comparing it to robust-constrained variants of PPO-Lagrangian (Ray et al.,, 2019) and RMCPMD (Wang et al.,, 2024), and find significant improvements in the penalised return in worst-case and average-case test performance on dynamics in the uncertainty set.
Related work
@@ 2.1 Robust constrained objectives
Russel et al., (2020) proposed the robust constrained policy gradient (RCPG) algorithm, which solves the inner problem based on either the robust cost or the robust constraint-cost, which ignores at least one part of the robustness problem in RCMDPs. Solving the RCMDP objective directly, the R3C objective (Mankowitz et al.,, 2020) combines the worst-case costs and worst-case constraint-costs by running distinct simulators with different parametrisations, each being run for one step from the current state. This unfortunately cannot be realised within a single simulator and is only applicable for discrete parametrisations. Additionally, it is also computationally expensive, even more so when there are multiple constraints. Alternatively, the robust Lagrangian objective trades off the cost and different constraint-cost according to whether the constraints are active while requiring only a single simulator with differentiable parameters (Bossens,, 2024).
In this paper, we use the robust Lagrangian objective with multiple constraints and rather than performing a single-step transition kernel update, we perform multi-step updates to the transition kernel at each macro-iteration, in a double-loop manner, to ensure that a near-optimal solution to the inner problem has been found, which in turn allows to provide guarantees on the near-optimality of the policy in solving the robust Lagrangian objective. Assuming that there exists at least one policy that is strictly feasible for all the constraints and all dynamics in the uncertainty set, the policy that optimises the robust Lagrangian will satisfy the worst-case constraint-costs and will, due to complementary slackness, optimise the worst-case value.
@@ 2.2 Adversarial transition kernels
Techniques for the design of adversarial transition kernels are of general interest for reinforcement learning. They represent different trade-offs between realism, scalability, computational cost, and sample efficiency. In the context of RMDPs and RCMDPs, they are used in the inner problem, which computes the worst-case transition kernel for the current policy. In general, convex uncertainty and rectangular uncertainty sets are required to make the problem tractable, since non-convex sets and non-rectangular set are in general NP Hard (Wiesemann et al.,, 2013).
@@
The scalability of these techniques is a key challenge. Linear programming based techniques (Russel et al.,, 2020) and techniques based on a small subset of simulators (Mankowitz et al.,, 2020) are not scalable. Instead, gradient-based methods can potentially be performed at scale. In RMDPs, gradient descent techniques for computing the inner problem have been proposed. Unfortunately, previous techniques do not have known guarantees (Abdullah et al.,, 2019; Bossens,, 2024) or provide results for non-rectangular uncertainty sets (which are NP Hard) (Li et al.,, 2023). Recent work has proposed transition mirror ascent (TMA), which comes with an $\mathcal{O}(1/T)$ convergence rate in the oracle-based setting under convex rectangular uncertainty sets, when using fixed learning rates, and $\tilde{\mathcal{O}}(e^{-T})$ convergence rates when using geometric schedules for the learning rate (Wang et al.,, 2024). A similar guarantee was earlier provided with a projected subgradient ascent algorithm, under slightly less general conditions, i.e. when the uncertainty set is a rectangular ball (Kumar et al.,, 2023).
@@
In RCMDPs, the Adversarial RCPG algorithm (Bossens,, 2024) provides a neural representation for transition kernels. Unfortunately, the algorithm does not have known theoretical guarantees. Our work takes note of the oracle-based convergence guarantees of the TMA algorithm of Wang et al., (2024), and extends the approach to design an algorithm for approximate TMA with $\mathcal{O}(1/T^{1/2})$ convergence rate in the sample-based setting. The result corresponds to the same oracle-based rate $\mathcal{O}(e^{-T})$ as in TMA (Wang et al.,, 2024) but accounts for the additional sample complexity from the estimation of the Q-values.
Defining the uncertainty set is an important challenge in the context of RMDPs and RCMDPs. The $\ell_{1}$ norm over the probability space is arguably the most widely used uncertainty set in the RMDP literature. Many other sets, such as $\ell_{p}$ sets, total variation distances, Chi-squared sets, and KL-divergences can be directly related to the $\ell_{1}$ norm. In turn, the bounds of the above sets can be derived from the true data based on relatively simple statistics. However, it is of great interest to construct tighter uncertainty sets with $\ell_{p}$ norms (e.g. Behzadian et al., (2019)) since tighter bounds imply a higher performing policy can be obtained in the true environment (as long as the set still includes the true environment). Beyond their relatively simple statistical derivation, $\ell_{p}$ sets are also fairly easy to project back onto the feasible set. Alternatives include models of uncertainty, suitable for indirect parametrisation and compatible with continuous state-action spaces, for instance the double sampling uncertainty sets, integral probability metrics uncertainty sets, and Wasserstein uncertainty sets (Zhou et al.,, 2023; Abdullah et al.,, 2019; Hou et al.,, 2007). Similarly, $\ell_{p}$ sets can also be applied to indirect parametrisations, which is compatible with the TMA algorithm (Wang et al.,, 2024). While in our main sample-based theorem, directly parametrised $\ell_{p}$ uncertainty sets are used to bound the distance from the nominal, the Lagrangian TMA and Approximate TMA methods can be applied with indirect parametrisations, such as the Gaussian Mixture PTK and the Entropy PTK (Wang et al.,, 2024), as also supported by our experiments.
@@ 2.3 Policy optimisation algorithms
Different policy opimisation algorithms have been proposed for RCMDPs and the related frameworks of CMDPs and RMDPs. As summarised in Table 1, the algorithms proposed in the literature work in the primal space, dual space, or a combination thereof, and they provide different guarantees under varying parametrisations.
In the CMDP literature on policy optimisation, the best convergence rates are obtained by PMD-PD (Liu et al., 2021b, ) in both the sample-based and oracle-based setting. Unlike other techniques using regularisation (Cen et al.,, 2022), the algorithm provides unbiased convergence results (i.e. they are expressed as the solution to the unregularised Lagrangian). The technique does not account for variability in the transition kernel, and therefore we extend it to the RCMDP setting.
@@
In the RCMDP literature, policy optimisers either come with no policy convergence guarantees (Russel et al.,, 2020; Mankowitz et al.,, 2020; Bossens,, 2024), basic feasibility and improvement results (Sun et al.,, 2024), or sub-optimal convergence points (Russel et al.,, 2023; Wang et al.,, 2022). Our sample-based theory provides results on par with CMDPs, namely an $\tilde{\mathcal{O}}(1/T^{1/3})$ convergence rate to the robust Lagrangian optimum, a result which is on par with the state-of-the-art PMD-PD algorithm (Liu et al., 2021b, ).
@@
Our algorithm is also practical in the sense that it can be implemented with deep RL techniques. In particular, we use Mirror Descent Policy Optimisation (MDPO) (Tomar et al.,, 2022) as the backbone of our algorithm’s implementation. The technique has been previously studied in traditional MDPs as a highly performing alternative to trust region based deep policy optimisers (e.g. PPO, Schulman et al., (2015) and TRPO, Schulman et al., (2017)). While the clipping implemented in PPO also allows a KL-constrained property, it often loses this property since due to averaging effects the policy is updated even if its clipping ratio is out of bound for some state-action pairs. TRPO computes the Fisher information, which can be expensive, and performs a KL constraint with a reversed order (i.e. $D_{\text{KL}}(\pi_{k},\pi)$ instead of $D_{\text{KL}}(\pi,\pi_{k})$ as used in mirror descent). The MDPO backbone not only leads to a novel algorithm for RCMDPs but also one for CMDPs, which turns out to be highly performant as an alternative to PPO-Lagrangian (Ray et al.,, 2019).
@@
Table 1: Overview of theoretical guarantees for policy optimisation algorithms in CMDPs and RCMDPs. Param indicates the parametrisation. Oracle rate indicates the convergence rate in the oracle setting. Sample rate indicates the convergence rate in the sample-based setting. The notation $\tilde{\mathcal{O}}$ hides logarithmic factors from the big O notation.
@@
Algorithm Algorithm type Param Convergence point Oracle rate Sample rate
@@
CPO (Achiam et al.,, 2017) trust region dual program with backtracking general feasible improvement / /
@@
RCPO (Tessler et al.,, 2019) primal-dual with projection general local CMDP optimum / /
@@
PCPO (Yang et al.,, 2020) trust region with projection general local CMDP optimum / /
@@
NPG-PD (Ding et al.,, 2020) primal-dual natural gradient ascent with projected subgradient descent softmax $\mathcal{O}(1/T^{1/2})$ $\tilde{\mathcal{O}}(1/T^{1/4})$
general $\mathcal{O}(1/T^{1/2})$ cost
$\mathcal{O}(1/T^{1/4})$ constraint
$\tilde{\mathcal{O}}(1/T^{1/4})$ cost
$\tilde{\mathcal{O}}(1/T^{1/6})$ constraint
@@
CRPO (Xu et al.,, 2021) primal softmax global CMDP optimum $\mathcal{O}(1/T^{1/2})$ $\tilde{\mathcal{O}}(1/T^{1/4})$
@@
PMD-PD (Liu et al., 2021b, ) primal-dual mirror descent softmax global CMDP optimum $\tilde{\mathcal{O}}(1/T)$ /
Sample-based PMD-PD (Liu et al., 2021b, ) primal-dual mirror descent softmax global CMDP optimum / $\tilde{\mathcal{O}}(1/T^{1/3})$
@@
Lyapunov-based RCPG (Russel et al.,, 2023) primal-dual, linear programming for inner problem general local RCMDP optimum / /

### v4

@@ Abstract
Safety is an essential requirement for reinforcement learning systems. The newly emerging framework of robust constrained Markov decision processes allows learning policies that satisfy long-term constraints while providing guarantees under epistemic uncertainty. This paper presents mirror descent policy optimisation for robust constrained Markov decision processes, making use of policy gradient techniques to optimise both the policy (as a maximiser) and the transition kernel (as an adversarial minimiser) on the Lagrangian representing a constrained Markov decision process. Our proposed algorithm obtains an $\tilde{\mathcal{O}}\left(1/T^{1/3}\right)$ convergence rate in the sample-based robust constrained Markov decision process setting. The paper also contributes an algorithm for approximate gradient descent in the space of transition kernels, which is of independent interest for designing adversarial environments in general Markov decision processes. Experiments confirm the benefits of mirror descent policy optimisation in constrained and unconstrained optimisation, and significant improvements are observed in robustness tests when compared to baseline policy optimisation algorithms.
1 Introduction
@@
Safe RL techniques have been primarily proposed within the constrained Markov decision process (CMDP) formalism (Altman, 1999). In infinite CMDPs, which are characterised by large or continuous state and action spaces, traditional solution techniques include linear programming and dynamic programming with Lagrangian relaxation. Following their empirical performance in high-dimensional control tasks (Lillicrap et al., 2015; Mnih et al., 2016), policy gradient algorithms have become a particularly popular option for safe RL in control applications (Achiam et al., 2017; Yang et al., 2020; Ding et al., 2020; Liu et al., 2021b; Xu et al., 2021; Paternain et al., 2023; Ying et al., 2023; Wu et al., 2024).
@@
The robustness to mismatches in training and test transition dynamics has been the main subject of interest in the robust MDP (RMDP) framework (Iyengar, 2005; Nilim & Ghaoui, 2005), which formulates robust policies by training them on the worst-case dynamics model within a plausible set called the uncertainty set. The framework has led to a variety of theoretically sound policy gradient approaches (Ho et al., 2021; Zhang et al., 2021; Kumar et al., 2023; Kuang et al., 2022; Wang et al., 2024; Li et al., 2023; Zhou et al., 2023; Wang et al., 2023). Recently, the approach has also branched out into the robust constrained MDP (RCMDP) framework (Russel et al., 2020; 2023; Wang et al., 2022; Bossens, 2024; Zhang et al., 2024; Sun et al., 2024), which seeks to provide robustness guarantees to CMDPs.
@@
The framework of RCMDPs is newly emerging and hence the theoretical development lags behind compared to CMDPs and RMDPs. At present, a few challenges have hampered the progress of RCMDP policy optimisation. First, since the framework is based on a worst-case cost function and one or more additional worst-case constraint-cost functions, it is challenging to implement in a single simulator since each of the robust cost functions may turn out to be based on a different transition kernel. This has lead to practical objectives that do not meet the original objective (Russel et al., 2020) or which only work for a discrete set of transition kernels (Mankowitz et al., 2020). Second, the state-of-the-art CMDP techniques for policy optimisation and RMDP techniques for adversarially optimising the transition dynamics have not yet widely made their way into the RCMDP literature, which has lead to a lack of theoretical guarantees. In RMDPs, state-of-the-art results are obtained for mirror descent based policy gradient algorithm (Wang et al., 2024), which applies mirror descent in the policy space and mirror ascent in the transition kernel space. In CMDPs, the policy mirror descent primal-dual (PMD-PD) algorithm is a related mirror descent based policy gradient algorithm which provides $\mathcal{O}(\log(T)/T)$ convergence rate is obtained in the oracle-based setting and an $\tilde{\mathcal{O}}(1/T^{1/3})$ convergence rate in the sample-based setting (Liu et al., 2021b). We then pose the question: using mirror descent based policy optimisation, is it intrinsically more difficult to optimise RCMDPs, or is it possible to obtain similar convergence rates?
@@
With the aim of solving these challenges and providing an affirmative answer to the above question, we draw inspiration from the robust Lagrangian objective (Bossens, 2024) and the success of mirror descent optimisation in RMDPs and CMDPs (Wang et al., 2024; Liu et al., 2021b). We show that it is indeed possible to obtain convergence rates similar to Sample-based PMD-PD in the sample-based RCMDP setting. For a theoretical analysis, we contribute Robust Sample-based PMD-PD, which extends the theory of Liu et al. (2021b) and Wang et al. (2024) to the RCMDP setting, while focusing on the softmax parametrisation within the sample-based setting. Doing so, we achieve the $\tilde{\mathcal{O}}(1/T^{1/3})$ convergence rate as also obtained by (non-robust) Sample-based PMD-PD. We then also extend the analysis to non-rectangular RCMDPs, albeit with a weaker sample-based convergence rate. For a practical implementation, we formulate MDPO-Robust-Lagrangian, which introduces a robust Lagrangian to Mirror Descent Policy Optimisation (MDPO) (Tomar et al., 2022). We evaluate MDPO-Robust-Lagrangian empirically by comparing it to robust-constrained variants of PPO-Lagrangian (Ray et al., 2019) and RMCPMD (Wang et al., 2024), and find significant improvements in the penalised return in worst-case and average-case test performance on dynamics in the uncertainty set.
In addition to techniques drawn from Liu et al. (2021b), Wang et al. (2024), and Tomar et al. (2022), our work designs novel techniques for theoretical analysis and contributes tools and results for practical experiments. In theoretical development, our work designs the following techniques:
  • Bounding the error due to sub-optimal dual variables: the proof of Theorem 4.2 of Wang et al. (2024) provides a result for bounding the robust regret in RMDPs by adding the tolerance of TMA and the regret on the current transition kernel. To extend this reasoning to RCMDPs, we provide an upper bound on the robust Lagrangian regret (Lemma 8) which in addition to these two terms also considers the suboptimality of the dual variables. This additional error is then shown to vanish in an average regret sense due to the constraint-costs averaging to $\mathcal{O}(\epsilon)$ across iterations, where $\epsilon>0$ is the error tolerance (Lemma 29).
  • Bounding errors due to transition kernel changes: a challenge in applying the telescoping analysis of Liu et al. (2021b) to RCMDPs is that some of the terms no longer cancel since the transition kernel changes after the policy update. To ensure that this analysis is still possible, Lemma 24 and Lemma 25 bound the error introduced by changes in the transition kernel using different telescoping sums.
  • Approximate TMA: previously, the TMA algorithm of Wang et al. (2024) relied on oracle information. To extend the algorithm to the sample-based setting, we have designed the Approximate TMA algorithm, which approximates the value function to perform mirror descent on transition kernels, and we derive its sample complexity as $\tilde{\mathcal{O}}(1/\epsilon^{2})$.
  • Continuous pseudo-KL divergence of occupancy: we present a pseudo-KL divergence for continuous state-action spaces and show it is a Bregman divergence, making PMD-PD techniques applicable to continuous state-action spaces.
  • Non-rectangular RCMDPs: while Wang et al. (2024) provides an analysis of TMA only for $s$-rectangular RMDPs and derives oracle-based convergence rates for the policy optimiser, our work applies the Appproximate TMA algorithm to non-rectangular RCMDPs and then derives a full sample-based analysis of the full algorithm optimising all three parameters $(\pi,p,\lambda)$.
From an experimental perspective, we highlight the following novelties:
  • Comprehensive ablation study: our study compares MDPO, PPO, and MCPMD algorithms for MDPs, CMDPs, RMDPs, and RCMDPs across three problem domains, including tests on the uncertainty set where the return and different penalised returns are evaluated.
  • Scalable implementation: the practical algorithm is highly scalable, making use of function approximators and other deep RL tools, as opposed to the pure Monte Carlo considered in Wang et al. (2024).
  • Learning rate schedules: we also present an empirical comparison of learning rate schedules, including traditional schedules fixed an linear schedules (Tomar et al., 2022), geometric schedules (Xiao, 2022; Wang et al., 2024), and a newly designed restart schedule. We show that for RCMDPs, the best performance is obtained by the MDPO-Robust-Augmented-Lag variant under batch-based restart schedules.
2 Related work
@@ 2.1 Robust constrained objectives
Russel et al. (2020) proposed the robust constrained policy gradient (RCPG) algorithm, which solves the inner problem based on either the robust cost or the robust constraint-cost, which ignores at least one part of the robustness problem in RCMDPs. Solving the RCMDP objective directly, the R3C objective (Mankowitz et al., 2020) combines the worst-case costs and worst-case constraint-costs by running distinct simulators with different parametrisations, each being run for one step from the current state. This unfortunately cannot be realised within a single simulator and is only applicable for discrete parametrisations. Additionally, it is also computationally expensive, even more so when there are multiple constraints. Alternatively, the robust Lagrangian objective trades off the cost and different constraint-cost according to whether the constraints are active while requiring only a single simulator with differentiable parameters (Bossens, 2024).
In this paper, we use the robust Lagrangian objective with multiple constraints and rather than performing a single-step transition kernel update, we perform multi-step updates to the transition kernel at each macro-iteration, in a double-loop manner, to ensure that a near-optimal solution to the inner problem has been found, which in turn allows to provide guarantees on the near-optimality of the policy in solving the robust Lagrangian objective. Assuming that there exists at least one policy that is strictly feasible for all the constraints under all dynamics in the uncertainty set, the policy that optimises the robust Lagrangian will satisfy the worst-case constraint-costs and will, due to complementary slackness, optimise the worst-case value.
@@ 2.2 Adversarial transition kernels
Techniques for the design of adversarial transition kernels are of general interest for reinforcement learning. They represent different trade-offs between realism, scalability, computational cost, and sample efficiency. In the context of RMDPs and RCMDPs, they are used in the inner problem, which computes the worst-case transition kernel for the current policy. In general, convex uncertainty and rectangular uncertainty sets are required to make the problem tractable, since non-convex sets and non-rectangular set are in general NP-Hard (Wiesemann et al., 2013).
@@
The scalability of these techniques is a key challenge. Linear programming based techniques (Russel et al., 2020) and techniques based on a small subset of simulators (Mankowitz et al., 2020) are not scalable. Instead, gradient-based methods can potentially be performed at scale. In RMDPs, gradient descent techniques for computing the inner problem have been proposed. Unfortunately, many previous techniques do not have known guarantees (Abdullah et al., 2019; Bossens, 2024). Recent work has proposed transition mirror ascent (TMA), which comes with an $\mathcal{O}(1/T)$ convergence rate in the oracle-based setting under convex rectangular uncertainty sets when using fixed learning rates, and $\mathcal{O}((1-\frac{1-\gamma}{M})^{T})$ convergence rates when using geometric schedules for the learning rate (Wang et al., 2024). A similar guarantee was earlier provided with a projected subgradient ascent algorithm, under slightly less general conditions, i.e. when the uncertainty set is a rectangular ball (Kumar et al., 2023).
@@
In RCMDPs, the Adversarial RCPG algorithm (Bossens, 2024) provides a neural representation for transition kernels. Unfortunately, the algorithm does not have known theoretical guarantees. Our work takes note of the oracle-based convergence guarantees of the TMA algorithm of Wang et al. (2024), and extends the approach to design an algorithm for approximate TMA with $\mathcal{O}(1/T^{1/2})$ convergence rate in the sample-based setting. The result corresponds to a slightly improved oracle-based rate $\mathcal{O}((1-\frac{1}{M})^{T})$ compared to the analysis of TMA in Wang et al. (2024) and further provides the additional sample complexity and estimation error from the estimation of the Q-values.
Despite the NP-Hardness result, recent work has also explored non-rectangular sets albeit with much weaker convergence guarantees (Li et al., 2023; Kumar et al., 2025). In addition to a non-tractable but globally optimal solution, Li et al. (2023) provides a conservative policy iteration algorithm for transition kernels that converges to a sub-optimal solution of an RMDP, with an error increasing based on the degree of non-rectangularity. While the rates in the paper hide the oracle for the value or gradient computation, it is the first to provide an approximate solution to general non-rectangular uncertainty sets. We exploit a similar technique and note its relation to the TMA algorithm, and then further analyse the overall algorithm with policy optimisation in the context of the non-rectangular RCMDP setting.
Obtaining tight and statistically based uncertainty sets is important for RMDP and RCMDP algorithms. The $\ell_{1}$ norm over the probability space is arguably the most widely used uncertainty set in the RMDP literature. For our purposes, it is convenient to use and arises naturally in our regret bounds. First, many other distance measures and statistics can be directly related to the $\ell_{1}$ norm, including $\ell_{q}$, total variation distances, Chi-squared sets, student-t statistics, Hoeffding or Chebyshev inequalities, the total variation distance, and KL-divergences. Beyond their statistical derivation and simplicity, $\ell_{q}$ sets are also fairly easy to project back onto the feasible set. The KL divergence is another distance metric with known analytical form for the transition kernel (see Table 7 in AppendixA) and with a direct connection to the norm. Second, as shown in Lemma 24, general uncertainty sets can be related to the $\ell_{1}$ norm, which appears naturally in the cumulative regret computation, and in this sense our main theorem is flexible to different divergences. In our convergence rate analysis, $\ell_{q}$ sets with small uncertainty budgets can potentially reduce one of the constants, although otherwise leading to the same convergence rate as in general uncertainty sets. Specialised algorithms have also been designed for the construction of tight weighted $\ell_{q}$ uncertainty sets while maintaining Bayesian confidence guarantees (e.g. Behzadian et al. (2019)).
Alternative models of uncertainty, suitable for indirect parametrisation and compatible with continuous state-action spaces, include examples such as the double sampling uncertainty sets, integral probability metrics uncertainty sets, and Wasserstein uncertainty sets (Zhou et al., 2023; Abdullah et al., 2019; Hou et al., 2007). Similarly, $\ell_{q}$ sets can also be applied to indirect parametrisations, which is compatible with the TMA algorithm (Wang et al., 2024). As supported by our experiments, Lagrangian TMA methods can similarly be applied with indirect parametrisations, such as the Gaussian Mixture PTK and the Entropy PTK (Wang et al., 2024).
@@ 2.3 Policy optimisation algorithms
Different policy opimisation algorithms have been proposed for RCMDPs and the related framework of CMDPs. As summarised in Table 1, the algorithms proposed in the literature work in the primal space, dual space, or a combination thereof, and they provide different guarantees under varying parametrisations.
In the CMDP literature on policy optimisation, many of the algorithms do not come with formal guarantees apart from local optima or improvement in the feasible space (Achiam et al., 2017; Tessler et al., 2019; Yang et al., 2020). Ding et al. (2020) present a primal-dual technique, NPG-PD, which obtains an $\mathcal{O}(1/T^{1/2})$ rate in the oracle-based setting for the softmax parametrisation, which extrapolates to an $\tilde{\mathcal{O}}(1/T^{1/4})$ bound in the sample-based setting. The work also presents a rate for a general parametrisation although it is considerably worse for the constraint-cost. CRPO applies NPG in the primal space with backtracking steps to ensure the constraint-cost remains satisfied, and is able to achieve comparable rates to NPG-PD (Xu et al., 2021). State-of-the-art results of $\tilde{\mathcal{O}}(1/T^{1/3})$ in the sample-based setting and $\tilde{\mathcal{O}}(1/T)$ in the oracle-based setting are obtained by PMD-PD, which applies a primal-dual mirror descent algorithm specifically with KL-divergence as a Bregman divergence (Liu et al., 2021b). Unlike other techniques using regularisation (Cen et al., 2022), the algorithm provides unbiased convergence results (i.e. they are expressed as the solution to the unregularised Lagrangian). PMD-PD does not account for variability in the transition kernel, and therefore we extend it to the RCMDP setting, obtaining comparable rates in the sample-based setting.
From the RMDP literature, the techniques by Wang et al. (2024) apply to the oracle-based rectangular RMDP setting, where for some $b\in(0,1)$ they achieve a state-of-the-art $\mathcal{O}(b^{T})$ convergence to the global optimum under a direct parametrisation of the policy and the transition kernel. For the softmax policy parametrisation, they report an $\mathcal{O}(1/T)$ oracle-based convergence rate in Theorem 4.5, a theorem which formulates the regret on the RMDP of interest based on the regret of an MDP optimiser passing through the same policy at the same iteration and with the transition kernel fixed to the nominal transition kernel at that iteration. However, their supporting Lemma B.5 incorrectly relies on a specific, favourable policy initialisation, which cannot be done in practice, rather than a general policy initialisation. The lemma also holds for any policy parameters at any iteration, such that no inference can be made about the regret of a particular policy in the RMDP optimisation simply by virtue of passing through the same parameters. Li et al. (2023) provides an $\mathcal{O}(1/T^{1/4})$ oracle-based result for non-rectangular RMDPs, the first convergence result for the global optimum of non-rectangular RMDPs considering the full algorithm solving both policy optimisation and the inner problem, while only requiring an approximate oracle for the gradient. The challenge of non-rectangularity shows up not only in the added complexity but also in additional error based on the degree of non-rectangularity. These works do not account for constraints or the samples to estimate the gradient; our work derives convergence results for the more challenging sample-based rectangular and non-rectangular RCMDP settings while maintaining similar global optimality properties.
@@
In the RCMDP literature, policy optimisers either come with no policy convergence guarantees (Russel et al., 2020; Mankowitz et al., 2020; Bossens, 2024), basic feasibility and improvement results (Sun et al., 2024), or sub-optimal convergence points (Russel et al., 2023; Wang et al., 2022). Our sample-based theory for rectangular uncertainty sets provides results on par with CMDPs, namely an $\tilde{\mathcal{O}}(1/T^{1/3})$ convergence rate to the robust Lagrangian optimum, a result which is on par with the state-of-the-art PMD-PD algorithm (Liu et al., 2021b). This compares favourably to the $\mathcal{O}(1/T^{1/4})$ oracle-based rate of Wang et al. (2022) which comes only with local convergence guarantees. We also provide a sample-based theory for non-rectangular RCMDPs, where an $\tilde{\mathcal{O}}(1/T^{1/5})$ rate is achieved. Both of our mentioned convergence rates account for all the samples required, including those for gradient estimation, transition kernel updates, policy updates, and dual variable updates.
@@
Our algorithm is also practical in the sense that it can be implemented with deep RL techniques. In particular, we use MDPO (Tomar et al., 2022) as the backbone of our algorithm’s implementation. The technique has been previously studied in traditional MDPs as a highly performing alternative to trust region based deep policy optimisers (e.g. PPO, Schulman et al. (2015), and TRPO, Schulman et al. (2017)). While the clipping implemented in PPO also allows a KL-constrained property, it often loses this property since due to averaging effects the policy is updated even if its clipping ratio is out of bound for some state-action pairs. TRPO computes the Fisher information, which can be expensive, and performs a KL constraint with a reversed order (i.e. $D_{\text{KL}}(\pi_{k},\pi)$ instead of $D_{\text{KL}}(\pi,\pi_{k})$ as used in mirror descent). The MDPO backbone leads to a novel algorithm for RCMDPs as well as for CMDPs, which turns out to be highly performant as an alternative to PPO-Lagrangian (Ray et al., 2019).
@@
Table 1: Overview of theoretical guarantees for policy optimisation algorithms in CMDPs and RCMDPs. MDP indicates the type of MDP, with prefixes C, R, and RC having their obvious meanings, while rR indicates rectangular robust and nR indicates non-rectangular robust. Param indicates the parametrisation. Oracle rate indicates the convergence rate in the oracle setting. Sample rate indicates the convergence rate in the sample-based setting. Other notations: $\tilde{\mathcal{O}}$ hides logarithmic factors from the big O notation; $b\in(0,1)$.
@@
Algorithm Algorithm type MDP Param Convergence point Oracle rate Sample rate
@@
(Achiam et al., 2017) trust region dual program with backtracking C general feasible improvement / /
@@
(Tessler et al., 2019) primal-dual with projection C general local optimum / /
@@
(Yang et al., 2020) trust region with projection C general local optimum / /
@@
(Ding et al., 2020) primal-dual NPG with projected subgradient descent C softmax global optimum $\mathcal{O}(1/T^{1/2})$ $\tilde{\mathcal{O}}(1/T^{1/4})$
@@
(Xu et al., 2021) primal unconstrained NPG with backtracking C softmax global optimum $\mathcal{O}(1/T^{1/2})$ $\tilde{\mathcal{O}}(1/T^{1/4})$
@@
(Liu et al., 2021b) primal-dual mirror descent C softmax global optimum $\tilde{\mathcal{O}}(1/T)$ $\tilde{\mathcal{O}}(1/T^{1/3})$
(Wang et al., 2024) double-loop mirror descent/ascent rR direct global optimum $\mathcal{O}(b^{T})$ /
(Li et al., 2023) projected descent and CPI ascent nR direct global with error $\delta_{\mathcal{P}}$ $\mathcal{O}(1/T^{1/4})$ /
@@
(Russel et al., 2023) primal-dual, linear programming for inner problem rRC general local optimum / /